OpenStudy (anonymous):

Five eighths of a number x is 3 more than one half of another number y. what does this equation look like. (FAN AND MEDAL QUESTION)

OpenStudy (anonymous):

I need to solve for y

OpenStudy (anonymous):

(a)y = (5x/8) + 3/2 (b)y = 110 + 2x (c)y = (5x/4) + 6 (d)y = (5x/4) - 6

OpenStudy (anonymous):

I think maybe a